OpenStudy (anonymous):

\[z=2^{-(x^2+y^2)}\] solve for y

OpenStudy (campbell_st):

take the base 2 log of both sides \[\log _{2} = -(x^2 + y^2)\] \[y^2 = -1(\log _{2}z + x^2)\] then take the square root of both sides to find y

OpenStudy (campbell_st):

oops missing a z on the left hand side of my 1st statement should be \[\log _{2} z = -(x^2 + y^2)\]